PHOTODISSOCATION DYNAMICS OF THE S_{1} AND S_{2} STATES OF CH_{3}ONO: ENERGY DISTRIBUTION IN THE CH_{3}O PHOTOFRAGMENT

CHRISTOPHER J. ANNESLEY, \underline{ANDREW E. BERKE}, F. FLEMING CRIM, Department of Chemistry, University of Wisconsin-Madison, Madison, Wisconsin 53706.

We studied the photodissocation dynamics of methyl nitrite excited to the S_{1} and S_{2} states by rediscovering the resonance enhanced multiphoton ionization of the CH_{3}O photofragment. While the NO photofragment’s state distibutions are well known, the state distributions in the methoxy had not previously been determined.